OpenStudy (anonymous):

i am horrible at math can anyone help? this is scientific notation btw medical X-rays with a wavelength of about 10 (-10)m, can penetrate completely throughout the skin Q- Ultraviolet rays which cause sun burns only penetrate top level of skin it has a wavelength about 1,000 times the wavelength of an X-ray, find the wavelength of ultraviolet rays. show your work -_- im horrible at math... does anybody understand this its like I'm reading greek or something

OpenStudy (anonymous):

dang thick chick looks like you got one hard problem on your hands

OpenStudy (anonymous):

i know -_- this is depressing

OpenStudy (anonymous):

well y don't u just talk to me and take a brake

OpenStudy (anonymous):

do you understand any of it?

OpenStudy (anonymous):

no what i would do is not do it and ask your teacher about it tomorrow in class

OpenStudy (anonymous):

lol i do online and the teacher is just as stupid as i am he makes us do textbook :/

OpenStudy (aum):

Wavelength of X-rays = \(\large 1.0 \times 10^{-10} ~ m \) Wavelength of Ultraviolet rays = 1000 times the wavelength of X-rays = \(\large 1000 \times 1.0 \times 10^{-10} ~ m = 1.0 \times 10^3 \times 10^{-10} = 1.0 \times 10^{3 - 10} = \\ \large 1.0 \times 10^{-7} ~ m\)

OpenStudy (anonymous):

and that is his problem he is making you do it out of a text book

OpenStudy (anonymous):

omg @aum your a genius ^.^ jesus has answered my prayers (hallelujah!!)

OpenStudy (aum):

lol. You are welcome.

OpenStudy (anonymous):

lol aum is smart as hell

OpenStudy (anonymous):

can you help with another?? @aum

OpenStudy (anonymous):

what is it

OpenStudy (aum):

I can try one more here. After that close this question and post a new one.

OpenStudy (anonymous):

ok this is the question the wavelengths of visible light are between 3.8 X 10 (-7)m and 7.6 X 10(-7)m Are these wavelengths longer or shorter than those of ultraviolet rays? explain.

OpenStudy (anonymous):

aum were u at she needs your help

OpenStudy (anonymous):

shhhh let the genius concentrate @williamauthorlee

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

i wonder how he understands it o.O im not even a blonde and half the crap it said made me feel like i was reading a foreign language

OpenStudy (anonymous):

that question made me feel stuid

OpenStudy (aum):

Wavelength of Visible light: \(\large 3.8 \times 10^{-7} \text{ to } 7.6 \times 10^{-7} \text{ m }\) Wavelength of Ultraviolet rays: \(\large 1.0 \times 10^{-7} \text{ m }\) Since \(\large 1.0 \times 10^{-7} \lt 3.8 \times 10^{-7} \), the wavelengths of Visible light is LONGER than the wavelength of Ultraviolet rays.

OpenStudy (anonymous):

you make me feel dumb -_- how do u understand this @aum

OpenStudy (aum):

Lots of practice and interest in math. Different things come easy to different people. I can't sing or dance or play musical instruments or write poetry or paint or ...... millions of things.

OpenStudy (anonymous):

I'm alright with math but this lesson just threw me off... big time but thanks this helped a lot

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

dang how do u understand this

OpenStudy (anonymous):

lol he practices and has better understanding with math than we do and I'm sure he's more advanced

OpenStudy (anonymous):

me to

OpenStudy (anonymous):

so thugchick do u have a email

OpenStudy (anonymous):

lol i do but i can't log into it

OpenStudy (anonymous):

dang y